# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v5

@@ 3.1 Datasets
Meetroom dataset [27] is another dataset for dynamic view synthesis that contains 3 scenes captured by 13 synchronized cameras at $1280\times 720$ resolution. The scenes are 300 frames long and captured at 30 FPS.
To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$ as shown in Fig. 2. Similar to those in the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS. The cameras are arranged in two rows and placed in an arc facing the scenes. The camera in the center of the upper row is designated for validation. We synchronize the multi-view videos and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation. Our dataset features more challenging dynamics, including significant non-rigid human motions and transient effects like screen slides, compared to the DyNeRF dataset. To comprehensively assess the scalability of our method, we provide 5 scenes with a duration exceeding 200 seconds (6000 frames), which is over 10 times longer than the previous DyNeRF dataset. Our dataset is collected to evaluate the scalability of our method and to support further research in the community. We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes. While the faces in the videos may contain identifiable information, this information will be used solely for research and will not be disclosed to any third party. All videos were collected with the explicit consent of all participants.
@@ 3.2 Implementation details
Our implementation is based on HexPlane [7], Instant-NGP [31], and NGP-pl [25]. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on both datasets. Following most of the state-of-the-art methods [28, 27, 17, 45, 43], we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We follow [27] and use $1280\times 720$ resolution on the Meetroom dataset. We use $1280\times 720$ resolution for our dataset to fit into the host memory. We use the importance sampling approach described in [28] at a batch size of 1024 rays. The model parameters are optimized by Adam [23], with hyper parameters $\epsilon=10^{-15},\beta=(0.9,0.96)$ for stability. The learning rate is set as 0.001 and scheduled by cosine function. We set the hash table sizes as $2^{P_{1}}=2^{19}$ in the first branch and $2^{P_{2}}=2^{14}$ in the subsequent branches. The structural parameters $(L,F)$ in the spatial hash encoders are set as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for $\eta_{init}=18k$ iterations, and for each of the subsequent branches we train for $\eta_{aux}=3k$ iterations using the previous branch for initialization for the DyNeRF dataset and our dataset. We use $\eta_{init}=33k$ and $\eta_{aux}=3k$ for the Meetroom dataset. We use LeakyReLU activation in the MLP regressors. Strategies for updating the occupancy grids are identical with [31]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision. The episode length $T_{episode}$ is set as $30$ only for videos longer than 300 frames in our own dataset. We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene.
@@ 3.3 Results and scalability analysis
We compare our method with both online and offline baselines. As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [48], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms the two additional baselines in the continual learning setting. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from the parameter isolation like StreamRF [27] and INV [48], our method achieves fast convergence and compact model size simultaneously. It leverages the hash tables to ensure fast convergence and reuses the features from differently configured spatial hash tables to ensure high scalability. Compared with prevailing offline methods [28, 34, 7, 17, 45, 55], our method only requires caching $T_{chunk}=10$ frames into the buffer by continual learning and thus reduces memory occupation by $85\%$. Our method also reduces $40\%$ less model size (to $0.4$MB/frame) by leveraging the feature fusion of the base and the auxiliary branches and delivers comparable quality to the offline methods.
@@
Figure 4: The results of the proposed CD-NGP on DyNeRF dataset.
@@
Table 2 and Fig. 5 show the quantitative and qualitative comparisons on the meetroom dataset. Fig. 6(a) is the PSNR curve of the compared methods on the trimming scene. We use the released configuration of StreamRF [27] for the meetroom dataset to obtain the results. As shown in Table 2, HexPlane-GR and HexPlane-PR produce significantly lower quality than the other methods on the meetroom dataset, we thus exclude them from Fig. 6(a) for better visualization. The metrics in Table 2 and the curves in Fig. 6(a) show the advantage of our method over both online and offline baselines. It outperforms the implicit MLP-based online method [48] regarding reconstruction quality, speed, and model size. Moreover, our method consumes only a fraction of storage than the explicit voxel-based online baseline [27] while maintaining comparable reconstruction quality as compared in Fig. 5. When compared with offline baselines, our method produces comparable quality to MixVoxels [45] and outperforms HexPlane [7] in terms of quality and speed by much less memory cost.
Figure 6: Combined PSNR curves for various scenes and methods. (a) The trimming scene in the Meetroom dataset. Solid and dotted lines denote offline and online methods respectively. (b) The first 1200 frames of seminar scene in our dataset. The solid lines and dotted lines denote offline and online methods respectively. (c) The first 2400 frames of seminar scene in our long video dataset. (d) the sear steak scene in the DyNeRF dataset. Solid and dotted lines denote symmetric and asymmetric setting, respectively.
@@
We compare the results of different spatial representations on the DyNeRF dataset and list the parameter numbers in Table 3. The alternatives of spatial representations include CD-MERF and CD-Plane. The element-wise product for plane representation in [17] produces unsatisfactory results in the experiments, we thus remove the $P_{m}(x,y,z)$ in Eq. 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation. As shown in Table 3, the hybrid representation CD-MERF achieves high compression and comparable model sizes to methods for static scenarios [31, 21] ($\sim$60MB) without significant quality loss. We also adopt the 3D+4D representation similar to that in [34] for comparison. This representation is implemented by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq. 5, i.e. $\Gamma(x,t)=\bigoplus_{m=1}^{L}P_{m}(x,y,z,t)$. However, the asymmetric hash sizes (using $(P_{1},P_{2})=(19,14)$) in the 3D+4D representation results in serious feature collisions and deliver bad results (PSNR < 16 dB on 3 scenes). We thus set $(P_{1},P_{2})=(14,14)$ for the 3D+4D representation in Table 3 to keep the size of the model close to our CD-NGP for fair comparison. The metrics show the advantage of our separated encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs. 3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
@@
As introduced in Section 2.3, our method shows superior scalability: the total size of the model is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, where $N_{chunk}$ is the number of chunks in the sequence. To prove the scalability of our method, we provide quantitative comparisons on our proposed long multi-view video dataset in Table 4 and Fig. 6(b) to further show the superior scalability of our method. We split the sequence into 300-frame chunks for the compared methods since the offline baseline [7] is designed for 300-frame sequences in the DyNeRF dataset. We set the hash table size as $2^{12}$ in the auxiliary branches and use $T_{chunk}=5$ in our model. As shown in Table 4, our method reduces the memory and storage consumption by at least 4×4$\times$4 × with comparable quality to HexPlane [7]. Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences. This is because the majority of the parameters in our method are used for the auxiliary branches, and the base branch only takes up a small fraction of the parameters. Consequently, applying smaller hash tables in the auxiliary branches and leverage the re-use of the features in the base branch leads to a significant reduction in the model size.
@@ 3.4 Ablation studies
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the lengths of the hash tables of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 5. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the auxiliary branches. The whole model contains hash encoders, MLP regressors, and the occupancy gridd, the sizes of the models are thus larger than the value deduced from the combinations of the hash table and MLP regressors. Although larger hash tables lead to less hash collisions, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the superiority of the asymmetric hash table design and prove the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 6(d). Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The $(18,18)$ in Table 5 configuration delivers the best results (30.55 dB in PSNR) on the DyNeRF dataset, but at the cost of an exceptionally large model size (638MB). However, the CD-NGP-C configuration with $(P_{1},P_{2})=(19,14)$ listed in Table 3 achieves a comparable PSNR (30.52 dB) with a much smaller model size (114MB, $80\%$ compression). The advantage comes from our re-use of the spatial features encoded by different branches: the base branch with a large hash table recovers more environment details, and the auxiliary branches could thus use fewer parameters to cope with recent changes only. The combination of the base encode and the auxiliary encoders lead to a more efficient and scalable model.
@@
Our continual learning pipeline is also compatible with a variety of representations for the time axis. Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [36] and MLP with frequency-encoding [16]. We replace the temporal hash encoding in our method with the ones mentioned above on DyNeRF dataset and show the results in Table 6. Our time latent code slightly outperforms others on DyNeRF dataset. The robustness against changes in the temporal representations also shows the viability of the continual learning method.
@@
Full composition in Table 7 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 14 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Table 7 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 14 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 9, the compared composition methods in Figs. 9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.

### v6

@@ 3.1 Datasets
The Meetroom dataset [27] is another dataset for dynamic NVS that recorded by 13 synchronized cameras with $1280\times 720$ resolution at 30 FPS. Each of the 3 scenes contains 300 frames.
Although existing datasets have been widely used for dynamic NVS, their limited lengths (300 frames) prevent the evaluation of the scalability of the methods on long video sequences. To further validate the effectiveness of our method on long video sequences, we introduce a new dataset comprising 5 dynamic scenes at a resolution of $3840\times 2160$ as shown in Fig. 2. Similar to those in the DyNeRF dataset [28], these scenes are captured by 20 GoPro cameras operating at 30 FPS. The cameras are arranged in two rows and placed in an arc facing the scenes. The camera in the center of the upper row is designated for validation. We synchronize the multi-view videos and incorporate chessboard images within the scenes to facilitate camera calibration and pose estimation. Our dataset features more challenging dynamics, including significant non-rigid human motions and transient effects like screen slides, compared to the DyNeRF dataset. To comprehensively assess the scalability of our method, we provide 5 scenes with a duration of 6000 frames. We hope this dataset will serve as a valuable benchmark for future research in dynamic novel view synthesis, especially in long video sequences. We have carefully reviewed the videos to ensure they do not contain any controversial or offensive content before using them for research purposes. While the faces in the videos may contain identifiable information, this information will be used solely for research and will not be disclosed to any third party. All videos were collected with the explicit consent of all participants.
@@ 3.2 Implementation details
We adopt the hash-encoding and tiny MLP implementation of tiny-cuda-nn [33] by Instant-NGP [34], and customize the CUDA-based training and rendering pipelines for the dynamic scenes based on [25]. The data preparation process is based on HexPlane [7] and LLFF [31]. Please see the supplementary material for more detailed explanations and analysis. All the metrics are measured on a single RTX 3090 GPU if not otherwise specified. Similar to the previous work [7, 28], we select the view in the center with index 00 for evaluation on all datasets. Following most of the state-of-the-art methods [28, 27, 17, 50, 48], we train and evaluate the model at resolution $1352\times 1014$ on the DyNeRF dataset. We follow [27] and use $1280\times 720$ resolution on the Meetroom dataset. We use $1280\times 720$ resolution for our dataset to fit into the host memory. We use the importance sampling approach described in [28] at a batch size of 1024 rays. The model parameters are optimized by Adam [23], with hyper parameters $\epsilon=10^{-15},\beta=(0.9,0.96)$ for stability. The learning rate is set as 0.001 and scheduled by a cosine function. We set the hash table sizes as $2^{P_{1}}=2^{19}$ in the first branch and $2^{P_{2}}=2^{14}$ in the subsequent branches. The structural parameters $(L,F)$ in the spatial hash encoders are set as $(12,2)$ if not otherwise specified. At the base branch, we initialize the models and train for $\eta_{init}=18k$ iterations, and for each of the subsequent branches we train for $\eta_{aux}=3k$ iterations using the previous branch for initialization for the DyNeRF dataset and our dataset. We use $\eta_{init}=33k$ and $\eta_{aux}=3k$ for the Meetroom dataset. We use LeakyReLU activation in the MLP regressors. Strategies for updating the occupancy grids are identical with [34]. To save storage, we use only one shared occupancy grid for all the branches. We report the model sizes with the occupancy grid under bfloat16 precision and parameters in float32 precision. The episode length $T_{episode}$ is set as $30$ only for videos longer than 300 frames in our dataset. We set the dimensions of the latent feature $\mathbf{H}$ in Eq. 3 as 48. We use MLPs with two hidden layers and the dimensions of the hidden layers in Eqs. 3 and 4 are set as 128, and 64 respectively. We use MLPs with identical structures for different branches. We set the resolutions of the hash encoders of the CD-NGP within $(16,2048)$ times the scale of the scene.
@@ 3.3 Results and scalability analysis
We compare our method with both online and offline baselines. Our method shows significantly better efficiency and quality than the other online methods. As demonstrated in Table 1, our CD-NGP is 32×32$\times$32 × faster than the MLP-based continual representation INV [53], $\times4$ smaller than the voxel-based continual representation StreamRF [27] and significantly outperforms these baselines in quality metrics. The undesirable results from HexPlane-PR and HexPlane-GR show the limitations of the regularization-based and replay-based methods in dynamic scenes: unlike the static scenes, the objects are not stationary across time and the models with limited complexity are not capable of reconstructing unlimited transient effects. Benefiting from parameter isolation like StreamRF [27] and INV [53], our method achieves fast convergence and compact model size simultaneously. It leverages the hash tables to ensure fast convergence and reuses the features from differently configured spatial hash tables to ensure high scalability.
Compared with most of the prevailing offline methods [28, 37, 7, 17, 50], our method only requires caching $T_{chunk}=10$ frames into the buffer by continual learning and thus reduces memory occupation significantly. Although the offline baseline 4DGS [61] consumes lower memory by using a lazy load strategy with a trade-off on training speed, their implementation consumes another 27 GB hard disk space on 300-frame DyNeRF dataset, and the space requirement also scales at a complexity of O($N_{frames}$) as the number of frames increases. On the contrary, our method leverages the continual learning strategy and enables directly decoding the frames from the videos without any additional storage consumption for training (scales at $O(1)$ complexity). It also balances training speed and model size by leveraging the feature fusion of the base and the auxiliary branches and delivers high quality.
@@
Figure 4: The results of the proposed CD-NGP on the DyNeRF dataset.
@@
Table 2 and Fig. 5 show the quantitative and qualitative comparisons on the Meetroom dataset. Fig. 6(a) is the PSNR curve of the compared methods on the trimming scene. We use the default setup of StreamRF [27] for the Meetroom dataset to obtain the results. As shown in Table 2, HexPlane-GR and HexPlane-PR produce significantly lower quality than the other methods on the Meetroom dataset, we thus exclude them from Fig. 6(a) for better visualization. The metrics in Table 2 and the curves in Fig. 6(a) show the advantage of our method over both online and offline baselines. We note that the Meetroom dataset provides only 13 views (much fewer than 20+ views in the DyNeRF dataset), which leads to limited coverage for multi-view stereo. As a result, COLMAP [44, 43] fails to generate high-quality dense point clouds, making it difficult for 4DGS [61] to obtain reliable initialization. Consequently, 4DGS does not produce valid results on this dataset and is excluded from the comparison. On the contrary, our method does not require initial point clouds and is more robust. It outperforms the implicit MLP-based online method [53] regarding reconstruction quality, speed, and model size. It obtains slightly lower metrics than the explicit voxel-based online method [27], which is likely due to the robustness advantage of explicit voxel representations under sparse-view conditions. When compared with offline baselines, our method produces comparable quality to MixVoxels [50] and outperforms HexPlane [7] in terms of quality and speed by much less memory cost.
Figure 6: Combined PSNR curves for various scenes and methods. (a) The trimming scene in the Meetroom dataset. Solid and dotted lines denote offline and online methods respectively. (b) The first 1200 frames of seminar scene in our dataset. The solid lines and dotted lines denote offline and online methods respectively. (c) The first 2400 frames of seminar scene in our long video dataset. (d) the sear steak scene in the DyNeRF dataset. Please note solid and dotted lines denote symmetric and asymmetric settings, respectively. The CD-NGP/MERF-$P_{1}$-$P_{2}$ annotation denotes different hash table sizes in the base and auxiliary branches defined in Section 2.3.
@@
We compare the results of different spatial representations on the DyNeRF dataset and list the parameter numbers in Table 3. The alternatives of spatial representations include CD-MERF and CD-Plane. The element-wise product for plane representation in [17] produces unsatisfactory results in the experiments. We thus remove the $P_{m}(x,y,z)$ in Eq. 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation. As shown in Table 3, the hybrid representation CD-MERF achieves high compression and comparable model sizes to methods for static scenarios [34, 21] ($\sim$60MB) without significant quality loss. We also adopt the 3D+4D representation similar to that in [37] for comparison. This representation is implemented by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq. 5, i.e. $\Gamma(x,t)=\bigoplus_{m=1}^{L}P_{m}(x,y,z,t)$. However, the asymmetric hash sizes (using $(P_{1},P_{2})=(19,14)$) in the 3D+4D representation results in serious feature collisions and deliver bad results (PSNR < 16 dB on 3 scenes). We thus set $(P_{1},P_{2})=(14,14)$ for the 3D+4D representation in Table 3 to keep the size of the model close to our CD-NGP for fair comparison. The metrics show the advantage of our separated encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs. 3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and leads to better results.
@@
As detailed in Section 2.3, our method exhibits superior scalability. The overall model size is $\mathcal{O}(N_{chunk}\cdot 2^{P_{2}}LF+2^{P_{1}}LF)$, where $N_{chunk}$ denotes the number of chunks in the sequence. We provide quantitative comparisons on our proposed long multi-view video dataset in Table 4 and Fig. 6(b). Following the default configurations on the DyNeRF dataset, the offline baselines HexPlane [7] and MixVoxels [50] are trained with a chunk length of $T_{chunk}=300$. Consistent with its DyNeRF configuration, 4DGS [61] adopts a lazy loading strategy, caching frames to disk in order to mitigate memory usage at the expense of training speed. In CD-NGP, we set the hash table size in the auxiliary branches to $2^{12}$ and adopt a compact chunk length of $T_{chunk}=5$. As shown in Table 4, our method reduces storage consumption by at least 4×4$\times$4 ×, delivers rendering quality comparable to HexPlane [7], and maintains low memory usage and fast training speed. Although 4DGS produces better results than our method, it consumes approximately 100 times more storage space. Compared with the results in Table 1, the performance gap in storage consumption is even larger in long video sequences. This is because the majority of the parameters in our method are used for the auxiliary branches, and the base branch only takes up a small fraction of the parameters. Consequently, applying smaller hash tables in the auxiliary branches and leveraging the reuse of the features in the base branch lead to a significant reduction in the model size.
@@ 3.4 Ablation studies
Benefiting from the multi-branch hash table design, our model is highly efficient and configurable. We change the lengths of the hash tables of the initial branch and the auxiliary branches and obtain the results on the DyNeRF dataset in Table 5. $(P_{1},P_{2})$ are the log2 values of the hash table lengths of the base branch and the auxiliary branches. The whole model contains hash encoders, MLP regressors, and the occupancy grid, the sizes of the models are thus larger than the value deduced from the combinations of the hash table and MLP regressors. Although larger hash tables lead to fewer hash collisions, the default configuration $(19,14)$ in CD-NGP produces better results than the naive $(16,16)$ configuration even if the storage consumption is $40.8\%$ lower. Moreover, the $(19,12)$ configuration of CD-NGP significantly outperforms the naive $(14,14)$ configuration at a similar storage cost. To further demonstrate the superiority of the asymmetric hash table design and prove the scalability of our method, we provide the PSNR curves of different hash table designs on the sear steak scene in Fig. 6(d). Equipped with a larger hash table in the base branch, the asymmetric setting $(19,14)$ and $(19,12)$ outperforms others in the beginning and produces comparable metrics to the naive $(16,16)$, $(14,14)$ configuration respectively in the subsequent chunks. The $(18,18)$ in Table 5 configuration delivers the best results (30.55 dB in PSNR) on the DyNeRF dataset, but at the cost of an exceptionally large model size (638MB). However, the CD-NGP-C configuration with $(P_{1},P_{2})=(19,14)$ listed in Table 3 achieves a comparable PSNR (30.52 dB) with a much smaller model size (114MB, $80\%$ compression). The advantage comes from our re-use of the spatial features encoded by different branches: the base branch with a large hash table recovers more environment details, and the auxiliary branches could thus use fewer parameters to cope with recent changes only. The combination of the base encoder and the auxiliary encoders leads to a more efficient and scalable model.
@@
Our continual learning pipeline is also compatible with a variety of representations for time. Currently, there are 2 other prevailing representations for the time axis alone: pure frequency encoding [39] and MLP with frequency-encoding [16]. We replace the temporal hash encoding in our method with them on the DyNeRF dataset and show the results in Table 6. Our time latent code slightly outperforms others on DyNeRF dataset. The robustness against changes in the temporal representations also shows the viability of the continual learning method.
@@
Full composition in Table 7 denotes the naive composition of two fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 14 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Table 7 denotes a naive composition of fields from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [50], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 14 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 9, the compared composition methods in Figs. 9(c) and 9(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results. Moreover, the feature combination in the latent space speeds up the inference by reducing the number of MLP queries. Please refer to the supplementary materials for more details.
@@
Table 10: Comparison between CD-NGP and MVSGS-FT on DyNeRF dataset.
[…]
CD-NGP 30.23 0.027 0.198
MVSGS-FT 26.05 0.086 0.258
@@
3.5 Further Discussion
CD-NGP leverages a highly scalable continual learning framework to tackle dynamic-scene NVS, with a principal benefit being a substantially reduced memory footprint (the feasible 14GB usage on the DyNeRF dataset). The multi-branch feature reusing design leads to fast convergence and high fidelity, and the encoder-decoder structure inherited from neural radiance fields (NeRF) [32] enables the model to represent the scenes compactly. CD-NGP supports varied kinds of encoders and field composition methods and exhibits high scalability in long video sequences. However, the volume rendering process is still required, which limits the real-time rendering capability of our method. Please see the supplementary materials for detailed latency analysis. Moreover, the current parameter isolation-based design does not support fusing features from different model branches, and thus the reconstruction quality of our method is lower than the state-of-the-art offline baselines. In addition, due to the hybrid hash encoder representations, our method suffers from more quality downgrades than the explicit baseline [27] in sparse-view scenarios.
We acknowledge that several recent cost volume-based multi-view stereo (MVS) NVS techniques also support real-time processing with low memory requirements and thus hold promise for dynamic-scene applications. To further evaluate the unique value of continual learning in this context, we include the state-of-the-art method MVSGS [30] as a representative baseline. MVSGS [30] leverages a pre-trained model to directly render novel views, as well as generating Gaussian point clouds to support per-scene fine-tuning for best quality. We refer to the direct-inference results as MVSGS-D and the fine-tuning results as MVSGS-FT. We note that the cost volume module of MVSGS does not support the standard $1352\times 1014$ resolution in the DyNeRF dataset, we thus use the default setup (including resolution) for MVSGS-D. Although processing the frames at $\sim 1$ FPS, MVSGS-D produces limited reconstruction quality on the DyNeRF dataset. We further compare MVSGS-FT and our CD-NGP in Table 10. We train the Gaussians for 5k iterations for each frame in the DyNeRF dataset. The training takes around 1.5 minutes for each frame, i.e. $\sim 8$ hours for the whole 300-frame sequence. The checkpoints consume around 60MB per frame (18 GB in total), which is significantly larger than the most recent explicit offline baselines for dynamic scenes. By comparing the quality metrics in Table 10 and the resource consumption, it is clear that our CD-NGP is more efficient: ours provides high-quality results, but consumes much less model size (113MB in total) and training time (75 minutes) by leveraging the re-use of the spatial features in the first branch. Hence, the significance of continual learning in dynamic-scene NVS is further validated. Please see the supplementary materials for details.